OpenStudy (anonymous):

What is the m∠ADC if angle DCB is 115?

HanAkoSolo (jamierox4ev3r):

What is the relationship between angle ADC and angle DCB?

OpenStudy (anonymous):

C:\Users\owner\Downloads\Garrett #9_files\triangle.gif

OpenStudy (anonymous):

answe choices 65, 115, 57.5,, and 90

OpenStudy (anonymous):

well there is no link i cant do anything

OpenStudy (anonymous):

copy and past my comment

OpenStudy (anonymous):

well its a file not a url

OpenStudy (anonymous):

bottom of the screen there should be attach file click it and find the file and post it

OpenStudy (anonymous):

the answer is the third one 57.5 because you can see that angle DCA is supplementary to angle DCB so 180 - 115 = 65 so because it is isocelese triangle it has to have to equal sides and two equal angles so the agle of CDA = 57.5 because 180 - 65 = 115 / 2 = 57.5

OpenStudy (anonymous):

does this help or youur still confused

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yw